# MIXPINN: Mixed-Material Simulations by Physics-Informed Neural Network

### v1

@@ I INTRODUCTION
Figure 1: MIXPINN pipeline concept. For a FEM simulation scene of a rigid medical device interacting with multi-anatomy structure, we transform it to GNN training on a physics-informed mesh-based graph under the influence of surface displacements. The model aims to predict the internal deformations by considering the interplay between soft tissue and multiple rigid components.
@@ III-B Input Features and Outputs
Considering a use case which involves applying an US probe onto a 3D anatomical model of the human torso where this mesh-based model contains positions of the mesh nodes, each labeled with 0,1,2,… to identify the anatomy membership (See IV-A for details), Table I summarizes the input features and outputs of the GAT model in this use case. Both node features denoted $\mathbf{X}$ and edge features $\mathbf{E}$ are used as inputs feeding into the GAT model. The 14-dimensional node features are constructed from 4 components. First, $[\delta x_{p},\delta y_{p},\delta z_{p}]$ is the displacement of the mesh’ surface points which contact the US probe directly, and the rest of the points have a value of $[0,0,0]$. Then, the initial position of all mesh nodes centered around the origin in both Cartesian coordinates $[x,y,z]$ and spherical coordinates $[r,\theta,\phi]$ are included. Finally, the rigid mask indicates the specific anatomy the node belongs to, in which 0 represents soft tissues and 1 to 5 represents vertebrae L1-L5 respectively. We include the one-hot encoding of the rigid mask $\mathbf{m}\in\{0,1\}^{5}$ in the node features. Edge features of dimension 6 include the length of the edge $l$, which corresponds to the Euclidean distance between the two endpoints of the edge. Moreover, edge features also contain the one-hot encoding of its anatomy membership $\mathbf{m}_{e}\in\{0,1\}^{5}$ similar to the node features. An edge is considered inside a specific rigid vertebra if both ends of the edge belong to that vertebra. Otherwise, it is considered as inside soft tissue with a label 0. The model output $\mathbf{Y}$ is the displacement of all mesh points in x, y and z directions, denoted $\delta x,\delta y,\delta z$.
@@ III-D Physics-informed Loss
where $|V|$ is the number of mesh nodes, $[x,y,z]$ is the ground truth displacements approximated by FEM and $[\hat{x},\hat{y},\hat{z}]$ is the predicted displacements from MIXPINN.
@@ IV-A Data Acquisition
Six segmentations from a human body Computed Tomography (CT) image are obtained from TotalSegmentator [27], namely, the main body trunk and lumbar vertebra L1 to L5. These images have $311\times 311\times 431$ 3D resolution with 1.5 mm voxel size. Using marching cubes, they are processed and combined to produce a multi-tissue triangular surface mesh for the body lumbar region. Then the surface mesh is used to generate the 3D tetrahedral mesh using The Computational Geometry Algorithms Library (CGAL) plugin in the SOFA framework with parameters cellRatio=2, cellSize=40, facetSize=30, facetAngle=30 and facetApproximation=1. Each mesh point and edge is labeled with the anatomy segment where 0 indicates soft tissue and 1 to 5 corresponds to vertebra L1-L5 respectively. Table II provides detailed statistics of the mesh generated.
@@
The mechanical properties of soft tissues are characterized by Stable Neo-Hookean material model [28], which is parametrized by a Poisson’s ratio of 0.45 and Young’s modulus of 25400 (Pa). Based on the significant stiffness discrepancy between soft tissues and spine, the lumbar vertebrae are modeled as rigid bodies [29]. This is achieved by subsetting the main body mesh and rigidifying the mesh nodes in each rigid vertebra separately. In this way, a multi-anatomy model is established. This mixed-material FE-model contains a deformable part and a rigid part which further divides to 5 sub-components. The US probe is also a rigid object in the simulation constructed from a Telemed L12 linear ultrasound transducer mesh [30]. To set up the boundary condition, some mesh nodes in the body anterior surface are fixed to support and stabilize the simulation. The initial probe positions are sampled from a structured mesh grid on the body back with 10 mm interval which produces 132 distinct positions. At each location, the US probe is positioned with four orientations corresponding to 0°, 45°, 90°, and 135° rotations. With each sampled pose, the probe moves at a uniform speed of 10 mm/s towards body back. We use the collision based contact response in SOFA with the friction coefficient set to 1 for a non-slippery contact. For each time step of 0.1 s, the quasi-static state of the system is resolved and computed. The displacements of all mesh points are exported. The simulation is ran for 20 time steps which corresponds to the probe compressed 2 cm into the body. In total 9796 experiment samples are created for training and validating the model.
@@ IV-B Held-out Experiment
To evaluate the model’s ability to generalize deformation predictions to unseen US probe pose, a specific testing strategy was used. The 132 simulation experiments of unique probe position are split by a ratio of $[0.7,0.2,0.1]$ so that their corresponding experiment data form the train, validation and test dataset. Table III reports the detailed statistics of the dataset created for GNN training. During training, a batch size of 4 is used and the data will be randomly shuffled after each epoch.
TABLE III: Statistics of dataset generated from FE-simulation used for GNN training
@@
Porbe Position 93 26 13 132
Porbe Pose 372 104 52 528
@@ V-D Runtime
TABLE VI: Average runtime to predict node displacements in one mesh deformation
@@
V-E Future Work
We plan to improve our method in a few possible areas. For example, the model can be modified to incorporate time series data to simulate beyond quasi-static scenario. So instead of considering each deformation separately, the model can take into account the displacement motion in the past time steps. This better align with real world applications or simulation, where the mesh deform gradually when interacting with continuous motion of external medical tools. On the other hand, the model performance can be further validated by conducting physical experiments.

### v2

@@ I INTRODUCTION
Figure 1: MIXPINN pipeline concept. For a FEM simulation scene of a rigid medical device interacting with multi-anatomy structure, we transform it to GNN training on a physics-informed mesh-based graph under the influence of surface contact displacements. The model aims to predict the internal deformations by considering the interplay between soft tissue and multiple rigid components.
@@ III-B Input Features and Outputs
Considering a use case which involves applying an US probe onto a 3D anatomical model of the human torso where this mesh-based model contains positions of the mesh nodes, each labeled with 0,1,2,… to identify the anatomy membership (See IV-A for details), Table I summarizes the input features and outputs of the GAT model in this use case. Both node features denoted $\mathbf{X}$ and edge features $\mathbf{E}$ are used as inputs feeding into the GAT model. The 14-dimensional node features are constructed from 4 components. First, $[\delta x_{p},\delta y_{p},\delta z_{p}]$ is the displacement of the mesh surface points which contact the US probe directly, and the rest of the points have a value of [0,0,0][0,0,0][ 0 , 0 , 0 ]. Then, the initial position of all mesh nodes centered around the origin in both Cartesian coordinates $[x,y,z]$ and spherical coordinates $[r,\theta,\phi]$ are included. Finally, the rigid mask indicates the specific anatomy the node belongs to, in which 0 represents soft tissues and 1 to 5 represents vertebrae L1-L5 respectively. We include the one-hot encoding of the rigid mask $\mathbf{m}\in\{0,1\}^{5}$ in the node features. Edge features of dimension 6 include the length of the edge $l$, which corresponds to the Euclidean distance between the two endpoints of the edge. Moreover, edge features also contain the one-hot encoding of its anatomy membership $\mathbf{m}_{e}\in\{0,1\}^{5}$ similar to the node features. An edge is considered inside a specific rigid vertebra if both ends of the edge belong to that vertebra. Otherwise, it is considered as inside soft tissue with a label 0. The model output $\mathbf{Y}$ is the displacement of all mesh points in x, y and z directions, denoted $\delta x,\delta y,\delta z$.
@@ III-D Physics-informed Loss
[…]
@@ IV-A Data Acquisition
Six segmentations from a human body Computed Tomography (CT) image are obtained from TotalSegmentator dataset [27], namely, the main body trunk and lumbar vertebra L1 to L5. These images have 311×311×431311$\times$ 311$\times$ 431311 × 311 × 431 3D resolution with 1.5 mm voxel size. Using marching cubes, they are processed and combined to produce a multi-tissue triangular surface mesh for the body lumbar region. Then the surface mesh is used to generate the 3D tetrahedral mesh using The Computational Geometry Algorithms Library (CGAL) plugin in the SOFA framework with parameters cellRatio=2, cellSize=40, facetSize=30, facetAngle=30 and facetApproximation=1. Each mesh point and edge is labeled with the anatomy segment where 0 indicates soft tissue and 1 to 5 corresponds to vertebra L1-L5 respectively. Table II provides detailed statistics of the mesh generated.
@@
The mechanical properties of soft tissues are characterized by the hyper-elastic Stable Neo-Hookean flesh model [28], which is parametrized by a Poisson’s ratio of 0.45 and Young’s modulus of 25400 (Pa) [29]. Based on the significant stiffness discrepancy between soft tissues and spine, the lumbar vertebrae are modeled as rigid bodies [30]. This is achieved by subsetting the main body mesh and rigidifying the mesh nodes in each rigid vertebra separately. In this way, a multi-anatomy model is established. This mixed-material FE-model contains a deformable part and a rigid part which further divides to 5 sub-components. The US probe is also a rigid object in the simulation constructed from a Telemed L12 linear ultrasound transducer mesh [31]. To set up the boundary condition, some mesh nodes in the body anterior surface are fixed to support and stabilize the simulation. The initial probe positions are sampled from a structured mesh grid on the body back with 10 mm interval which produces 132 distinct positions. At each location, the US probe is positioned with four orientations corresponding to 0°, 45°, 90°, and 135° rotations. With each sampled pose, the probe moves at a uniform speed of 10 mm/s towards body back. We use the collision based contact response in SOFA with the friction coefficient set to 1 for a non-slippery contact. The simulation is ran for 20 time steps which corresponds to the probe compressed 2 cm into the body. For each time step of 0.1 s, the quasi-static state of the system is resolved and computed. The displacements of all mesh points are exported. In this way, a total of 9796 experiment samples are created for training and validating the model.
@@ IV-B Held-out Experiment
To evaluate the model’s ability to generalize deformation predictions to unseen US probe position and pose, a specific testing strategy was used. The 132 simulation experiments of unique probe position are split by a train, validation, test ratio of [0.7,0.2,0.1][0.7,0.2,0.1][ 0.7 , 0.2 , 0.1 ]. Their respective probe poses (from 4 rotations) are also split accordingly. Table III reports the detailed statistics of the dataset created for GNN training. During training, a batch size of 4 is used and the data will be randomly shuffled after each epoch.
TABLE III: Statistics of the dataset used for GNN training generated from FE-simulation
@@
Probe Position 93 26 13 132
Probe Pose 372 104 52 528
@@ V-D Runtime
TABLE VI: Average runtime of one mesh deformation prediction
@@
V-E Discussion
By providing real-time yet accurate deformation prediction on multi-anatomy tissues, our method can be deployed to enhance intra-operative surgical planning or used for tool trajectory predictions. We plan to improve our method in a few possible areas. For example, the model can be modified to incorporate time series data to simulate beyond quasi-static scenario. On the other hand, the model performance can be further validated by conducting physical experiments on phantom or cadaver. The method can also be benchmarked on a wider range of clinical scenarios.
@@ VI CONCLUSIONS
ACKNOWLEDGMENTS
This work is part of the ”Learn to learn safely” project funded by a grant of the Hasler foundation (grant nr: 21039). Furthermore, we extend our sincere thanks to Hugo Talbot for his valuable consulting on the SOFA Framework.